how to solve equations representing pair of straight lines?.eg. 6xsquare + xy - ysquare=0 

Grade:11

1 Answers

SAGAR SINGH - IIT DELHI
13 years ago

6x2+xy-y2=0

6x2+3xy-2xy-y2=0

3x(2x+y)-y(2x+y)=0

(3x-y)(2x+y)=0

Now you get a pair of straight lines.

 

Please feel free to post as many doubts on our discussion forum as you can.

We are all IITians and here to help you in your IIT JEE preparation.

All the best